OpenStudy (anonymous):

Select an ordered pair from the choices below that is a solution to the following system of inequalities: y < -7 + 8x 8x + 3y > 24 (3, 10) (-1, -4) No solution (0, 0)

OpenStudy (anonymous):

Substitute the ordered pairs accordingly into your x and y and see which one will make the inequality true.

OpenStudy (anonymous):

so i think its is (3,10)

OpenStudy (anonymous):

y<-7+8x 8x+3y>24 10<-7+8(3) = 10<-7+24 = 10<17 TRUE 8(3)+3(10)>24 = 24+30>24 = 54>24 TRUE so its (3,10)